OpenStudy (anonymous):

Find the inverse of f(x) = 3x - 2

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

so it is x+2/2

OpenStudy (zzr0ck3r):

y = 3x-2 solve for x y+2 = 3x (y+2)/3 = x so the inverse is f(x) = (x+2)/3

OpenStudy (zzr0ck3r):

maybe that was a typo?

OpenStudy (anonymous):

yup